# Wake-Informed 3D Path Planning for Autonomous Underwater Vehicles Using A* and Neural Network Approximations

### v1

@@ 4.1 Path Planner Operation
The environment is first discretized into a three-dimensional grid $G\subset\mathbb{Z}^{3}$, where each cell (voxel) represents a discrete location in 3D space. The grid dimensions are defined to be ($[1,128],\ [1,128],\ [1,128]$) representing the 3D $128^{3}$ grid structure. Each individual cell in the grid corresponds to a node $n\subset G$. These nodes are connected to their neighbors based on a 26-connected grid. This considers movement in all directions: up, down, left, right, forward, backward and diagonals. For a node $n$, its set of neighbors $N(n)$ includes all adjacent nodes that are within the grid bounds and are traversable (not occupied by obstacles). The Euclidean distance (as denoted in Eq. (1)) is used to determine the distance between the current node n and the next node $n^{\prime}$.
@@
The cost of moving from node n to the next node n’ is defined based on the energy required to overcome drag forces $F_{D}$, in addition to the distance traveled $d$. This can be determined, as the velocity throughout the local environment is known across all cells, and an approximation of the force of drag ($F_{D}$) can be made via Eq. (2).
@@
in which $\rho$ denotes density, v denotes the relative velocity between the vehicle and the fluid domain, $C_{D}$ denotes the coefficient of drag and A denotes the cross-sectional area of the vehicle in the drag direction. In turn, $F_{D}$ is then used to approximate the energy ($E_{D}$) required to overcome drag over a distance, $d$, such that:
$E_{D}=F_{D}\times d$ (3)
Subsequently, $E_{D}$ is used to define the movement cost $c$ between two nodes $n$ and $n^{\prime}$ as seen below in $c(n,n^{\prime})$. A weighting term $\omega_{1}$ has also been introduced, which can be used to weight the importance of drag, which may be important, depending on the application.
$c(n,n^{\prime})=d(n,n^{\prime})\times\omega_{1}\cdot E_{D}(n^{\prime})$ (4)
in which $d(n,n^{\prime})$ is used to denote the Euclidean distance between nodes $n$ and $n^{\prime}$, and $E_{D}(n^{\prime})$ represents the energy cost due to drag at node $n^{\prime}$. This then results in the cost function comprising the following:
$g(n)=\sum_{k}^{i=1}c(n_{i-1},n_{i})=\sum_{k}^{i=1}d(n,n^{\prime})\times\omega_{1}E_{D}(n^{\prime})$ (5)
which can be further decomposed into the cost such that:
$g(n)=\sum_{k}^{i=1}d(n,n^{\prime})\times\omega_{1}\left[\frac{1}{2}\rho v^{2}C_{D}A\right]$ (6)
where $n_{0}$ is the starting node and $n_{k}$ is the final node, comprising the set of trajectory points such that $P=[n_{0},n_{1},\cdots,n_{k}]$. The initial starting location is $n_{0}$ which is randomly uniformly distributed on the rear entrance of the domain. The goal location, $n_{k}$, is localized to the center of the payload bay. The goal of the path planning optimization is to determine a path, $P^{*}$, from the starting location $n_{0}$ to the goal location $n_{k}$ that minimizes the total cost such that:
[…]
Two variants of this path planner are used, where the velocity (v) from the drag equation has different levels of fidelity. In the uninformed path planner, only the current field velocity ($v_{C}$) is known. In the informed path planner, the wake structure and current field ($v_{T}$) of the velocity is known. The $A^{*}$ algorithm is then used to plan the path through the grid $G\subset\mathbb{Z}^{3}$ in accordance with Eq. (7) between a starting node and goal node within a traversable graph space. The heuristic function estimates the cost to reach the goal from the starting node using the average drag energy per unit distance and the aforementioned Euclidean distance. This heuristic is defined to be the sum of the average drag energy per unit distance ($\overline{E_{D}}$) within the grid $G\subset\mathbb{Z}^{3}$ and distance such that:
$h(n)=\overline{E_{D}}\times d(n,n^{\prime})$ (8)
The heuristic is admissible as it will not overestimate the true minimum cost to reach the goal. This is ensured by the use of the average drag energy, which provides a reasonable estimate to guide the search within the domain, without comprising optimality. For each neighbor $n^{\prime}$ of the current node n, the algorithm calculates the tentative cost to reach $n^{\prime}$ via n and updates the scores accordingly. The nodes are stored in a priority queue order, sorted by their $f(n)=g(n)+h(n)$ values to ensure that the most promising nodes are explored first.
@@
To assess the importance of knowing the wake structures, two variants of the planner will be assessed. The first variant, $P^{*}_{C}$ will only be informed of the current within the computational domain. The second variant, $P^{*}_{W}$, will be informed of both the current and the vehicle’s wake structure. $P^{*}_{W}$ will be applied to the flow field across a range of wake fields behind a large underwater vehicle. The current and wake dataset was generated from prior work by the authors and is comprised of high fidelity computational fluid dynamics (CFD) data interpolated onto the discrete grid $G\subset\mathbb{Z}^{3}$ [1] [2]. The data contains a local current from 0.10 m/s to 5.00 m/s in increments of 0.10 m/s. Additionally, a strong propeller wake is included, with wake structures at 0° to 60° degrees of separation, simulating a turn of the vehicle in question. This was done via the simulation of an INSEAN E1619 propeller. The wake free velocity field ($v_{C}$) containing only the current, will be the same field, with the wake structure’s components filtered out and only global current retained.
@@ 4.2 Path Generation
For the wake structure dataset, with local current from 0.10 m/s to 5.00 m/s in increments of 0.10 m/s and angles of incidence from 0° to 60° degrees of separation in increments of 5°, there are a total of 500 distinct flow conditions. For each flow condition, 36 trajectories are generated with starting positions that are uniformly disturbed on the aft face of the domain, behind the XLUUV. An example of this is provided in Fig. 3. This image shows how the uniformly distributed start locations are initialized in the domain, before trajectory determination. This process occurs 36 times for each 500 distinct flow conditions, resulting in a total of 18,000 paths being generated for each planner. The same start positions are used for both the wake-informed and current-informed planners, allowing for direct comparison across relevant metrics.
Figure 3: Front side view of the domain with multiple trajectories plotted in red. Starting locations of the AUV are indicated in green. Larger vehicle geometry indicated in blue, contained in the wake structure of the body and propeller (teal).
@@ 4.3 Path Assessment Metrics
Energy capacity onboard AUVs remains a constant challenge, where small hull structures limit the physical size of battery systems, resulting in tightly constrained vehicles. It is important that the trajectories generated are assessed from an energy cost perspective, as well as the variance in energy cost, as these factors can have a significant effect on the viability of a given path. Additionally, when approaching a larger vessel, the wake structure shed by it can cause significant approach challenges, particularly with respect to effective control of the smaller vehicle [1] [2]. It is for this reason that the hydrodynamic effects should be assessed to determine the proportion of the trajectory where a control risk exists. We define this control risk to be both regions on the trajectory where there is a higher than expected flow speed, and where the velocity is fluctuating in the x, y or z direction.
To make this assessment, each generated path, $P$, is decomposed into a set of $N$ nodes such that:
@@
Each node $n_{i}$ in the set corresponds to a 3-dimensional coordinate $(x_{i},y_{i},z_{i})$ in the grid. The velocity at node $n_{i}$ is $V(n_{i})$. $\tilde{v}$ denotes the median velocity of the entire grid, and $\sigma_{v}$ denotes the standard deviation of the velocities in the grid. These values are used to identify high velocity cells, where we define the high velocity cell threshold $v_{\text{threshold}}$ to be a function of the velocity field’s standard deviation $\sigma_{v}$:
@@
After the velocity threshold for these cells is identified, the path $P$ is used to determine the number of cells along the trajectory that exceed this threshold. The number of high velocity cells is then summed over the trajectory, denoted by $N_{\text{high-velocity}}(P)$ and calculated via Eq. (11), where the cutoff threshold for a high velocity cell is denoted by the indicator function in Eq. (12) and given via $\delta_{\text{HV}}$.
$N_{\text{high-velocity}}(P)=\sum_{i=1}^{N}\delta_{\text{HV}}(V(n_{i}))$ (11)
$\delta_{\text{HV}}(v)=\begin{cases}1,&\text{if }v\geq v_{\text{threshold}}\\ 0,&\text{otherwise}\end{cases}$ (12)
To assess the fluctuations in velocity, or the turbulence along the path $P$, a similar method is applied. $D(n_{i})$ denotes the velocity disparity at a given node $n_{i}$. At any node where the disparity value is non zero, it is summed. This keeps track of the number of disparity cells along the path $P$, denoted by $N_{\text{disparity}}(P)$. This is determined via the application of Eq. (13). Again, the cutoff threshold for a velocity fluctuation cell is denoted by the indicator function $\delta_{\text{D}}$ defined as follows, using a non-zero offset $(\varepsilon=0.005)$ to ensure calculation robustness. This is provided in Eq. (14)
$N_{\text{disparity}}(P)=\sum_{i=1}^{N}\delta_{\text{D}}(D(n_{i}))$ (13)
$\delta_{\text{D}}(d)=\begin{cases}1,&\text{if }d\geq\varepsilon\\ 0,&\text{otherwise}\end{cases}$ (14)
To determine the energy cost of a path, the drag force $F_{D}(n_{i})$ at a given node $n_{i}$, is obtained from the cost array. Then, for each segment between nodes $n_{i}$ and $n_{i+1}$, the Euclidean distance $d_{i}$ is calculated via the application of Eq. (1) and the force of drag determined via Eq. (2). These are then combined and determined for each cell along the trajectory, allowing for a path energy, denoted by $E(P)$, calculation via Eq. (15).
$E(P)=\sum_{i=1}^{N-1}F_{D}(n_{i})\times d_{i}$ (15)
The total Euclidean distance (path length) along path $P$, denoted $L(P)$, is the sum of the distances between consecutive nodes where $d_{i}$ is as previously defined:
$L(P)=\sum_{i=1}^{N-1}d_{i}$ (16)
The path length in cells, denoted by $N_{\text{cells}}(P)$, is simply the number of nodes ($n$) on each given path:
$N_{\text{cells}}(P)=N$ (17)
@@ 5 Neural Network Design
Whilst A* calculation of a trajectory results in extremely optimal paths, this calculation requires more substantial computing resources than are typically available on an AUV within it’s limited computational hardware and energy budget, rendering it either impractical or impossible. Neural networks (NNs), in contrast, are capable of learning complex mappings from input features to outputs. This makes them ideal candidates to approximate the function of both the wake- and current-informed path planning A* algorithms [29]. Beneficially, NNs are significantly less computationally intensive (once trained) and offer the ability to rapidly generate and update the trajectories based on a new inputs, unlike the recalculations required for traditional A*. This section details the steps taken to develop, train and assess the NNs designed to approximate the two planners.

### v2

@@ 4.1 Path Planner Operation
The environment is first discretized into a three-dimensional grid $G\subset\mathbb{Z}^{3}$, where each cell (voxel) represents a discrete location in 3D space. The grid dimensions are defined to be ($[1,128],\ [1,128],\ [1,128]$) representing the 3D $128^{3}$ grid structure. Each individual cell in the grid corresponds to a node $n$. These nodes are connected to their neighbors based on a 26-connected grid, allowing movement to any adjacent voxel sharing a face, edge, or corner. For a node $n$, its set of neighbors $N(n)$ includes all adjacent nodes that are within the grid bounds and are traversable (not occupied by obstacles). The Euclidean distance between two nodes $n=(x_{n},y_{n},z_{n})$ and $n^{\prime}=(x_{n^{\prime}},y_{n^{\prime}},z_{n^{\prime}})$ is denoted by $d(n,n^{\prime})$ and is given by Eq. (1):
@@
The cost of moving from node $n$ to an adjacent node $n^{\prime}$ is defined based on the estimated energy required to overcome hydrodynamic drag force $F_{D}$ during that movement. This can be determined, as the velocity throughout the local environment is known (or estimated) within each cell of the grid $G$. The drag force ($F_{D}$) acting on the AUV is determined via Eq. (2). The AUV particulars and values used to enable accurate drag calculation for Eq. (2) are provided in B.
@@
In Eq. (2) $\rho$ denotes the fluid density, $v$ denotes the magnitude of the relative velocity between the vehicle and the fluid, $C_{D}$ denotes the coefficient of drag for the AUV, and $A$ denotes its cross-sectional area. The drag can be used in conjunction with the Euclidean distance to determine the energy cost associated with moving to $n^{\prime}$. To calculate this cost of traversal ($n$ to $n^{\prime}$), we evaluate the drag force using the fluid velocity conditions at the destination node $n^{\prime}$, denoted $F_{D}(n^{\prime})$. The subsequent energy $E_{D}(n,n^{\prime})$ required to overcome this drag, over the distance $d(n,n^{\prime})$, is approximated as by Eq. (3).
$E_{D}(n,n^{\prime})=F_{D}(n^{\prime})\times d(n,n^{\prime})$ (3)
where $F_{D}(n^{\prime})=\frac{1}{2}\rho v(n^{\prime})^{2}C_{D}A$, using the relative fluid velocity $v(n^{\prime})$ estimated at node $n^{\prime}$.
The energy cost $E_{D}(n,n^{\prime})$ is then used to define the movement cost $c(n,n^{\prime})$, also known as the edge weight, for traversing from node $n$ to node $n^{\prime}$. A weighting factor $\omega_{1}$ (typically set to 1 unless scaling is desired) can be included as denoted in Eq. (4) which states the movement cost as a function of the energy ($E_{D}$) which is comprised of both the fluid forces ($F_{D}$) and the movement distance ($d$).
$c(n,n^{\prime})=\omega_{1}\cdot E_{D}(n,n^{\prime})=\omega_{1}\cdot F_{D}(n^{\prime})\cdot d(n,n^{\prime})$ (4)
The A* algorithm searches for a path by maintaining the cost of the path found so far from the start node $n_{0}$ to the current node $n_{curr}$. This is known as the $g$-score, denoted by $g(n_{curr})$, and is calculated by summing the edge costs along the path given by $P=[n_{0},n_{1},\dots,n_{curr}]$. The equation for this is provided in Eq. (5).
$g(n_{curr})=\sum_{i=1}^{curr}c(n_{i-1},n_{i})$ (5)
The path planning problem is formulated as an optimization task. The design variable is the path $P$, represented as a sequence of connected nodes $[n_{0},n_{1},\dots,n_{k}]$ from a given start node $n_{0}$ to a specified goal node $n_{k}$. The objective is to find the optimal path $P^{*}$ that minimizes the total cumulative cost, $g(n_{k})$, required to reach the goal $n_{k}$, as stated by Eq. (6).
$P^{*}=\underset{P\in\mathcal{P}(n_{0},n_{k})}{\operatorname{arg\,min}}\ g(n_{k})=\underset{P=[n_{0},...,n_{k}]}{\operatorname{arg\,min}}\left(\sum_{i=1}^{k}c(n_{i-1},n_{i})\right)$ (6)
where $\mathcal{P}(n_{0},n_{k})$ is the set of all possible valid paths from $n_{0}$ to $n_{k}$.These valid paths are assessed via a heuristic function. The A* algorithm utilizes the heuristic function, $h(n_{curr})$, to estimate the minimum remaining cost from the current node $n_{curr}$ to the goal node $n_{k}$. This estimate guides the search efficiently towards the goal. In this work, the heuristic is defined as the product of the straight-line Euclidean distance $d(n_{curr},n_{k})$ between the current node ($n_{curr}$) and the goal ($n_{k}$). This is provided in Eq. (7).
$h(n_{curr})=\overline{E_{D}}\times d(n_{curr},n_{k})$ (7)
Here, $\overline{E_{D}}$ represents the global average of the energy term, which consists of the weighted drag force term, $\omega_{1}\cdot F_{D}(n^{\prime})$ (derived from Eq. 4). This heuristic combines the geometric distance to the goal with an overall estimate of the environmental energy resistance. Crucially, this heuristic is admissible because the average energy cost $\overline{E_{D}}$ will not overestimate the actual minimum energy cost per unit distance along any specific optimal path segment. Since the actual cost on the optimal path from $n$ to $n_{k}$ involves traversing specific cells with potentially lower-than-average or higher-than-average costs, using the global average to guide the search prevents systematic overestimation of the remaining cost. Admissibility is a required property for A* to guarantee finding the minimum cost path (Hart et al., 1968). While simple, this heuristic provides a computationally inexpensive way to focus the search, balancing distance and the energy landscape of the grid $G$.
The A* algorithm explores nodes based on the priority function $f(n)=g(n)+h(n)$. It maintains an ’open set’ (priority queue) of nodes to be evaluated, ordered by their $f$-score, and a ’closed set’ of nodes already evaluated. Starting with $n_{0}$, it iteratively selects the node $n$ with the lowest $f$-score from the open set, expands its neighbors $n^{\prime}$, calculates their tentative $g$-scores and $f$-scores, and updates them if a cheaper path is found. The process continues until the goal node $n_{k}$ is selected for expansion, at which point the optimal path is reconstructed by backtracking from the goal.
@@
To assess the importance of knowing the wake structures, two variants of the planner will be assessed. The first variant, $P^{*}_{C}$ will only be informed of the current within the computational domain. The second variant, $P^{*}_{W}$, will be informed of both the current and the vehicle’s wake structure. $P^{*}_{W}$ will be applied to the flow field across a range of wake fields behind a large underwater vehicle. The current and wake dataset was generated from prior work by the authors and is comprised of high fidelity computational fluid dynamics (CFD) data interpolated onto the discrete grid $G\subset\mathbb{Z}^{3}$ (Cooper-Baldock et al., 2024) (Cooper-Baldock et al., 2023). More details about this are provided in A. The data contains a local current from 0.10 m/s to 5.00 m/s in increments of 0.10 m/s. Additionally, a strong propeller wake is included, with wake structures at 0° to 60° degrees of separation, simulating a turn of the vehicle in question. This was done via the simulation of an INSEAN E1619 propeller. The wake free velocity field ($v_{C}$) containing only the current, will be the same field, with the wake structure’s components filtered out and only global current retained.
4.2 Graphical Depiction
Fig. 3 details the A* path planner methodology. The positions of the XLUUV and AUV (x, y, z) are used to discretize a 3D domain in the $128^{3}$ structure as discussed. This grid structure is then augmented with either the current or wake flow data as denoted (blue and purple respectively). Following this, the path planner is started, with the start and goal locations ($n_{0},n_{k}$) defined. A priority queue is initialized before the assessment process for the neighboring nodes ($n^{\prime}$) using the search heuristic ($h$) is undertaken. This process is iterated, with the best node selected based on the cost until the goal location is reached. At this point, the planner is stopped, as a trajectory has been generated.
Figure 3: Operation of the A* planner architecture. Domain is initialized with the position of the XLUUV and AUV relative to each other. Method of flow data initialization (current only or wake) indicated in blue and purple respectively. Operation of the A* planner node selection indicated in orange.
@@ 4.3 Path Generation
To compare the effectiveness of the wake-informed A* planner and the current-informed A* planner, many trajectories needed to be generated for assessment. This was achieved through the use of the CFD dataset (A). This dataset contained 500 distinct flow fields, each of which were characterized by a different speed and angle pairing, and resulted in a different wake flow field behind the XLUUV.
For each of these 500 instances, the location of the AUV was randomly generated on the rear boundary of the 3D domain (depicted in Fig. 4). If the current-informed planner was being assessed, the grid velocities $G(v)$ would be overwritten with the mean velocity value ($\overline{v}$) in all cells, resulting in a domain comprising of only the global current. If the wake-planner was being assessed, then the wake and detailed velocity structure of the CFD data would be retained.
The randomized start points of the AUV, of which 36 were generated for each of the 500 fields, were held consistent for both the current- and wake-informed planners. From these starting positions, the path planning process (outlined in Fig. 3) would be followed, until the AUV had been navigated to the center of the payload bay onboard the XLUUV (the goal $n_{k}$). Using 36 trajectories for each of the 500 flow conditions resulted in 18,000 paths being determined for the current- and wake-informed planners respectively.
An example instance, of 1 of the 500 distinct flow fields, is provided in Fig. 4. In this image, the 36 paths generated for the wake-informed A* planner are shown. The 36 starting locations of the AUV on the rear boundary are denoted by the green spheres. The trajectories towards and into the XLUUV are denoted by the red lines, and the wake velocity data ($v_{wake}$) is provided in blue, with the XLUUV located in the center of the wake field.
Figure 4: Example path generation of multiple A* trajectories shown. Each trajectory is indicated in red, with the goal of avoiding the wake flow structure (indicated in blue). Domain is shown from a front 3/4 side view. The starting location of each AUV trajectory is indicated by the green sphere. The trajectory goal location is the center of the payload bay within the XLUUV, including the final stage recovery, through the boundary layer. The XLUUV body geometry is contained within the blue wake flow.
@@ 4.4 Path Assessment Metrics
The energy capacity onboard an AUV is extremely limited. This is due to their small physical size and inability to house large battery systems (Cooper-Baldock et al., 2023). Due to the limited energy budgets, there are several important performance metrics that should be assessed for any trajectory.
A successful trajectory should be energy efficient, short in physical length and also minimize the traversal of adverse high velocity or turbulent regions of the domain $G$. To assess these metrics, several performance metrics have been designed. These are the method to determine the energy cost of the path ($E$), the length of the path ($L$), the number of high velocity cells traversed $N_{\text{high-velocity}}(P)$) and the number of turbulent cells traversed ($N_{\text{high-velocity}}(P)$).
To make these assessments, each generated path, $P$, is decomposed into a set of $N$ nodes such that:
@@
Each individual node ($n_{i}$) in the set corresponds to a 3-dimensional coordinate $(x_{i},y_{i},z_{i})$ in the grid. The velocity at node $n_{i}$ is $v(n_{i})$. $\tilde{v}$ denotes the median velocity of the entire grid $G$, and $\sigma_{v}$ denotes the standard deviation of the velocities in the grid. Crucially, whilst the current-informed planner operated without knowledge of the wake flow in $G$, it will be assessed in a domain containing the wake flow, in order to assess the impact of the wake on the metrics identified above and the impact of accounting for the wake in the trajectory planning stage. $\sigma_{v}$ is used to identify high velocity cells, where we define a high velocity cell to be 1 standard deviation or more from the median velocity ($\tilde{v}$). Where an individual node ($n_{i}$) encounters a cell that is more than 1 standard deviation from the velocity field’s mean value in the grid $G$, it is recorded, indicating the traversal of a high velocity cell. This threshold is given by Eq. 9
@@
The number of cells along the trajectory that exceed this threshold is denoted by Eq. 10. This equation tracks the number of high velocity cells via $N_{\text{high-velocity}}(P)$ and is calculated via Eq. (10), where the cutoff threshold for a high velocity cell is denoted by the indicator function in Eq. (11) and given via $\delta_{\text{V}}$ as mentioned.
$N_{\text{high-velocity}}(P)=\sum_{i=1}^{N}\delta_{\text{V}}(n_{i})$ (10)
$\delta_{\text{V}}(n_{i})=\begin{cases}1,&\text{if }v(n_{i})\geq v_{\text{threshold}}\\ 0,&\text{otherwise}\end{cases}$ (11)
To assess the degree of turbulence encountered over the trajectory, a similar method is applied. Where the velocity is found to change at a given node (via $v^{\prime}(n_{i})$), it is summed, and thus tracks the number of velocity fluctuations recorded over the path of the trajectory (through $N_{\text{turbulent}}(P)$). The CFD flow field data, determined via ANSYS Fluent, may have slight fluctuations due to numerical precision, which can result in small changes in velocity despite uniform freestream velocity conditions. To stop these cells from being incorrectly tracked, a small offset is applied $(\varepsilon=0.005)$ ensuring that the velocity fluctuation, sufficient to be recorded as a turbulent cell transition, must be more than 0.5% of $v$. The two equations to enable this tracking are Eq. (12) and (13) respectively.
$N_{\text{turbulent}}(P)=\sum_{i=1}^{N}\delta_{\text{T}}(n_{i})$ (12)
$\delta_{\text{T}}(n_{i})=\begin{cases}1,&\text{if }v^{\prime}(n_{i})\geq\varepsilon\\ 0,&\text{otherwise}\end{cases}$ (13)
To determine the energy cost of a path, the drag force ($F_{D}$) as denoted by Eq. 2 and distance ($d$) denoted by Eq. 1 of each node ($n_{i}$) and the next ($n_{i+1}$) are summed along the trajectory via the application of 14. This yields the energy cost of a path $E(P)$ in joules.
$E(P)=\sum_{i=1}^{N-1}F_{D}(n_{i},n_{i+1})\times d(n_{i},n_{i+1})$ (14)
The total path length ($L$) along the set of nodes in $P$, denoted $L(P)$, is the sum of the distances between consecutive nodes where $d_{i}$ is as previously defined by 1.
$L(P)=\sum_{i=1}^{N-1}d(n_{i},n_{i+1})$ (15)
@@ 5 Neural Network Design
While the A* algorithm, when using an admissible heuristic as employed here (see Section 4.1), is guaranteed to find the path with the minimum cumulative cost on the discretized grid graph (Hart et al., 1968), this calculation requires substantial computing resources. Such resources are typically unavailable on an AUV with it’s limited computational hardware and energy budget, rendering it either impractical or impossible to undertake such planning in real time.
Neural networks (NNs), in contrast, are capable of learning complex mappings from input features to outputs. This makes them ideal candidates to approximate the function of both the wake- and current-informed path planning A* algorithms (LeCun et al., 2015). Beneficially, NNs are significantly less computationally intensive (once trained) and offer the ability to rapidly generate and update the trajectories based on a new inputs, unlike the recalculations and computational resources required for traditional A*.
To investigate the potential for real-time path generation, two distinct neural network (NN) models were developed and trained. Each NN was specifically designed to approximate the output of one of the A* planners: one network learns from the trajectories generated by the current-informed A* ($C.I._{A*}$) and the other learns from the wake-informed A* ($W.I._{A*}$) trajectories. This separation allows each network to specialize in mimicking the behavior of its corresponding A* algorithm, which operates on different environmental input data in the domain $G$ (current-only vs. current and wake).